OpenStudy (anonymous):

lim as x approaches 0 from the right of x^(1/2)ln(7x). how do you know to change it to ln(7x)/x^(-1/2)?

OpenStudy (anonymous):

because just by making it negative is the same as putting it on the denominator of the fraction. X^1/2 = sqrt(x) and x^-1/2= 1/sqrt(x)... but in order to solve that problem, you need to graph it, look at the x=0 value from the right.

OpenStudy (anonymous):

ok well we were told to use lahopital rule and i cant understand which to put on top or bottom

OpenStudy (anonymous):

ah la hôpital, unfotunaly our AP calc class has not gotten that far, otherwise id love to help, sorry :(

OpenStudy (anonymous):

haha np lol ty though